OpenStudy (3psilon):

The angular acceleration α for some particle moving in a circle depends linearly on its angular velocity ω. α = A ω where A is some constant. The initial angular velocity is ω0. The initial angular position is ω0/A. What is θ(t)?

OpenStudy (anonymous):

θ(t) = ωo t + 1/2 α t^2

OpenStudy (anonymous):

\[\theta (t) = \frac{ \omega _{0} }{ A } +\omega _{0} t +\frac{ 1 }{ 2 } A \omega _{0}t^2\]